A medical researcher believes that a drug changes the body's temperature. Seven test subjects are randomly selected and the body temperature of each is measured. The subjects are then given the drug, and after 30 minutes, the body temperature of each is measured again. The results are listed in the table below. Is there enough evidence to conclude that the drug changes the body's temperature? Let d=(body temperature after taking drug)−(body temperature before taking drug). Use a significance level of α=0.01 for the test. Assume that the body temperatures are normally distributed for the population of people both before and after taking the drug. Subject 1 2 3 4 5 6 7 Temperature (before) 98 100.1 99 98.9 100.7 98.8 98.7 Temperature (after) 97.8 99.5 99.3 98.1 100.2 98.4 98.2 Step 3 of 5 : Compute the value of the test statistic. Round your answer to three decimal places.
Answer #1
Subject Temperature (Before) X Temperature (After) Y di=Y-X di-d_bar (di-d_bar)^2
1 98 97.8 -0.2 0.18 0.0324
2 100.1 99.5 -0.6 -0.22 0.0484
3 99 99.3 0.3 0.68 0.4624
4 98.9 98.1 -0.8 -0.42 0.1764
5 100.7 100.2 -0.5 -0.12 0.0144
6 98.8 98.4 -0.4 -0.02 0.0004
7 98.7 98.2 -0.5 -0.12 0.0144
SUM = -2.7 0.7488
MEAN = -0.38

AN Edia -2.7 in=7 D = J = Idi-2,7 -0.38] -0.38 Adid)2 = 0.7488 $ds {gli-4)? 0.7488 7-1 n- Sd= 0.35327 = 0.01 tarzyn+1 = 60.00ALSO USING MINITAB,

Paired T-Test and CI: Temperature (After) Y, Temperature (Before) X

Paired T for Temperature (After) Y - Temperature (Before) X

N Mean StDev SE Mean
Temperature (Aft 7 98.7857 0.8859 0.3348
Temperature (Bef 7 99.1714 0.9160 0.3462
Difference 7 -0.385714 0.353217 0.133503


99% CI for mean difference: (-0.880668, 0.109240)
T-Test of mean difference = 0 (vs not = 0): T-Value = -2.89 P-Value = 0.028

SINCE P- value =0.028 > Alpha =0.01

Hence, we fail to reject H0

PLZ RATE IT POSITIVE.